# Vanishing conductivity-like Gilbert damping in mirror-symmetric van der Waals ferromagnets

### v1

@@ Abstract
The enhancement of damping by nodal lines is confirmed by elevating $E_{F}$ in 1QL Fe3GaTe2. Two scattering rates, $\gamma=3$ meV and 20 meV, were selected, and the calculated damping is plotted in Fig. 4(c) as a function of energy. Near the original $E_{F}$, $\alpha$ is minimal and insensitive to energy shifts since the intraband transitions are forbidden and the nodal lines are far above $E_{F}$. For $E-E_{F}>200$ meV, as energy approaches the nodal lines, $\alpha$ increases rapidly for both $\gamma$ values due to significantly enhanced interband transitions. We further investigate the scattering rate dependence of $\alpha$ at various energies. As shown in Fig. 4(d), the typical monotonic resistivity-like damping is observed at $E-E_{F}=100$ and $200$ meV, where the nodal lines are not significantly involved. In these scenarios, the interband contribution exhibits a nearly linear dependence on $\gamma$ and tends to saturate at large $\gamma$. When scattering occurs at $E-E_{F}=300$ and $400$ meV, damping becomes substantial even at very low $\gamma$. This is because the interband transitions always occur at the crossing point, irrespective of broadening. Thus, the damping displays a very weak dependence on $\gamma$. It also implies that the common assumption $\alpha_{\rm interband}\propto\rho$ does not apply in these instances.
@@
The interband transitions also explain the higher $\alpha$ calculated for 3QL Fe3GaTe2 compared to 1QL, as shown in Fig 3(b). With mirror symmetry maintained, we analyze the band structure for 3QL [47]. The tripling of atoms in the unit cell results in a greater number of bands in 3QL compared to 1QL. Consequently, the denser bands in 3QL lead to a significantly larger band overlap under the same broadening $\gamma$, culminating in a substantially larger interband contribution to damping for 3QL than for 1QL.
@@
Figure 5: Calculated damping anisotropy between out-of-plane and in-plane magnetization for Fe3GaTe2 (a) and Fe3GeTe2 (b).
Discussion and conclusion.—Our analysis of the intraband and interband contributions to Gilbert damping reveals an inherent anisotropy in the mirror-symmetric vdW FM material. When the magnetization is oriented out-of-plane, the intraband contribution is symmetry-forbidden, thereby reducing the associated damping. However, by rotating the magnetization into the plane using external fields, the mirror symmetry is broken, leading to the recovery of the intraband (conductivity-like) damping. Consequently, the Gilbert damping can be controlled by adjusting the magnetization orientation without altering the scattering rate. This anisotropic damping is particularly significant in clean systems with relatively low scattering rates, as the intraband contribution is inversely proportional to $\gamma$. The calculated damping anisotropy between out-of-plane and in-plane magnetization for Fe3GaTe2 is depicted in Fig. 5(a). For both 1QL and 3QL, a substantial anisotropy ratio of up to $10^{4}\%$ is observed at small $\gamma$, and it exceeds 200% for the bulk material. This anisotropy decreases with increasing $\gamma$ and disappears in the strongly disordered limit, which is governed by interband transitions. The latter depends on the band overlap and is insensitive to symmetry constraint. Fe3GeTe2, which shares the same lattice structure and symmetry as Fe3GaTe2, exhibits the similar properties in its Gilbert damping [47]. Figure 5(b) presents the calculated damping anisotropy of Fe3GeTe2, which is slightly less than that of Fe3GaTe2. This difference arises because the interband contribution is relatively larger in Fe3GeTe2 due to the closer proximity of nodal lines to the Fermi level.
@@
In summary, our investigation into the Gilbert damping of 2D vdW ferromagnetic metals, Fe3GaTe2 and Fe3GeTe2, has uncovered the role of mirror symmetry in governing the damping behavior. The conductivity-like component of damping in 1QL vanishes under out-of-plane magnetization due to symmetry-prohibited intraband transitions. However, this symmetry can be broken either by tilting the magnetization or by varying the number of layers, which reinstates the intraband contribution and substantially enhances damping. Furthermore, symmetry-protected crossings between mirror-symmetric and antisymmetric bands create nodal lines where the interband transitions are pronounced, leading to a significant damping enhancement. Our findings not only elucidate the microscopic origins of the observed monotonic temperature-dependence of Gilbert damping but also predict an inherent damping anisotropy in these metals that awaits experimental validation. Additionally, we propose that voltage control offers a promising avenue for tuning the damping of Fe3GeTe2 and Fe3GaTe2 thin films by modulating the Fermi level relative to the nodal lines.
@@ Acknowledgements.
This work was supported by the National Natural Science Foundation of China (Grants No. 12374101 and No. 12174028) and the Open Project of Guangdong Provincial Key Laboratory of Magnetoelectric Physics and Devices, No. 2022B1212010008.

### v3

@@ Abstract
The enhancement of damping by nodal lines is confirmed by elevating $E_{F}$ in 1QL Fe3GaTe2. Two scattering rates, $\gamma=3$ meV and 20 meV, were selected, and the calculated damping is plotted in Fig. 4(c) as a function of energy. Near the original $E_{F}$, $\alpha$ is minimal and insensitive to energy shifts since the intraband transitions are forbidden and the nodal lines are far above $E_{F}$. For $E-E_{F}>200$ meV, as energy approaches the nodal lines, $\alpha$ increases rapidly for both $\gamma$ values due to significantly enhanced interband transitions. We further investigate the scattering rate dependence of $\alpha$ at various energies. As shown in Fig. 4(d), the typical monotonic resistivity-like damping is observed at $E-E_{F}=100$ and $200$ meV, where the nodal lines are not significantly involved. When scattering occurs at $E-E_{F}=300$ and $400$ meV, damping becomes substantial even at very low $\gamma$. This is because the interband transitions always occur at the crossing points, irrespective of broadening. Thus, the damping displays a very weak dependence on $\gamma$. It also implies that the common assumption $\alpha_{\rm interband}\propto\rho$ does not apply in these instances.
@@
The interband transitions also explain the higher $\alpha$ calculated for 3QL Fe3GaTe2 compared to 1QL, as shown in Fig 3(b). With mirror symmetry maintained, we analyze the band structure for 3QL (Sec. 3 in Supplemental Material [54]). The tripling of atoms in the unit cell results in a greater number of bands in 3QL compared to 1QL. Consequently, the denser bands lead to a larger band overlap under the same broadening $\gamma$, culminating in a substantially larger interband contribution to damping for 3QL than for 1QL.
Discussions and conclusions.—The predicted features in Gilbert damping of mirror-symmetric vdW ferromagnets can be directly measured in experiments. The symmetry-forbidden intraband transitions with out-of-plane magnetization results in a low damping. By rotating the magnetization using an external field, the mirror symmetry is broken, leading to the recovery of the intraband (conductivity-like) damping. Consequently, the Gilbert damping can be controlled by adjusting the magnetization orientation without altering the scattering rate, which is particularly significant in clean systems. The calculated damping anisotropy between out-of-plane and in-plane magnetization for Fe3GaTe2 is depicted in Fig. 5. For both 1QL and 3QL, a substantial anisotropy ratio exceeding $10^{4}\%$ is observed at small $\gamma$, and it exceeds 200% for bulk. This damping anisotropy is expected to be observable using FMR or time-resolved magneto-optical Kerr effect. Additionally, the monotonic (non-monotonic) temperature-dependent damping for odd (even) layers can also be experimentally measured, as suggested by Fig. 3(b).
@@
Figure 5: Calculated damping anisotropy between out-of-plane and in-plane magnetization for Fe3GaTe2.
@@
In summary, our investigation into the Gilbert damping of 2D vdW FM metals has uncovered the crucial role of mirror symmetry in governing damping behavior. In 1QL Fe3GaTe2, the conductivity-like damping vanishes under out-of-plane magnetization due to symmetry-forbidden intraband transitions. This results in an ultralow Gilbert damping at small scattering rate. However, this symmetry can be broken either by tilting the magnetization or by varying the number of layers, which reinstates the intraband contribution and substantially enhances damping. Furthermore, symmetry-protected crossings between mirror-symmetric and antisymmetric bands create nodal lines where the interband transitions are pronounced, leading to a significant damping enhancement. Voltage control offers a promising avenue for tuning the damping by shifting the Fermi level relative to these nodal lines. These nodal lines and associated topological electronic states may have interesting interactions with topological magnons in vdW FM systems [57, 58, 59, 60]. Our findings not only elucidate the microscopic origins of the observed monotonic temperature dependence of Gilbert damping but also predict a universal mechanism to achieve the ultralow Gilbert damping in 2D vdW FM materials, with no lower limit.
@@ Acknowledgements.
This work was supported by the National Key Research and Development Program of China (2024YFA1408500), the National Natural Science Foundation of China (Grants No. 12374101 and No. 12574115), the Open Fund of the State Key Laboratory of Spintronics Devices and Technologies (Grant No. SPL-2408) and the Open Project of Guangdong Provincial Key Laboratory of Magnetoelectric Physics and Devices (2022B1212010008).